# Dual-Model Distillation for Efficient Action Classification with Hybrid Edge-Cloud Solution

### v1

@@ 1 Introduction
Recent advances in Artificial Intelligence (AI) have shown that increasing model size leads to substantial performance gains across various domains, from language models (Dubey et al., 2024; Jiang et al., 2023) to video-language models (Lin et al., 2023; Wu et al., ). Large-scale models, with billions of parameters, consistently achieve state-of-the-art results on various benchmarks (Zhao et al., 2023). However, these models can be challenging to deploy when translating benchmark performance into real-world applications (Minaee et al., 2024) due to their high runtime costs.
@@
One critical area where this limitation becomes evident is in action classification for elderly care (Sawik et al., 2023). A task such as predicting falls through non-intrusive monitoring (Maldonado-Bascon et al., 2019; Elwaly et al., 2024) is a critical yet complex task that benefits from large-scale video-language models (Elwaly et al., 2024). Despite their high accuracy, these oversized models are not feasible for deployment on common elderly care appliances due to hardware constraints (Chen and Ran, 2019). To address this, we propose a hybrid edge-cloud solution where small models run locally on edge devices while larger models deployed in the cloud are accessed only when necessary. However, the core technical challenge here remains: how to effectively align the knowledge and capabilities of the small and large models to ensure seamless collaboration between them. While Model or Knowledge Distillation techniques (Hinton, 2015; Gou et al., 2021) have been extensively explored, their primary goal is to compress the knowledge of large models into smaller ones. Yet, this process often falls short in collaborative systems where both small and large models have unique strengths. The smaller model, after distillation, may not fully exploit its distinct capabilities or optimally collaborate with its larger counterpart (Narayan et al., 2022).
@@
To overcome these limitations, we introduce Dual-Model Distillation (DMD), a novel approach to enhance cooperation between small and large models. Unlike traditional distillation, which focuses on unidirectional knowledge transfer, DMD treats both models as teacher models to co-teach and generate synthetic data to train a student model (i.e., the switcher model). The DMD process eliminates the need for additional human-labeled data by leveraging the agreement or disagreement between the two teacher models’ zero-shot inferences to create training data. Through instruction-supervised finetuning (Chung et al., 2024), the switcher model learns to make optimal decisions about when to invoke the larger model. Experiments conducted on the dataset used in Fall Detection - Eldercare Robot datase Elwaly and Abdellatif (2023) demonstrate the effectiveness of our approach. Surprisingly, our pipeline not only demands less computational overhead than just using the large model, resulting in 39.5% cost savings, but also outperforms the isolated large model by 4.6% in terms of F1 score. Furthermore, we anticipate that the DMD concept will provide substantial opportunities for future research on knowledge integration across multiple deep learning models.
@@ 2 Methodology
As shown in Figure 1, our hybrid edge-cloud solution introduces a companion switcher model using the hidden features of the input data extracted from the small model, which serves as the foundation of its reasoning capacity. With the help of the switcher model, our edge model can not only predict the falling probability but also accurately estimate the probability that the large model’s result aligns with its own. By incorporating this switcher model, the edge model deliberates when to invoke the cloud model, thereby keeping system expenses under a pre-defined budget. In our case, a threshold of alignment probability, which yields optimal performance in predicting the falling probability during training, determines when to offload data to the large model during inference time. Regarding more complex decision-making scenarios, we expect that a more sophisticated utility-based threshold design Wang and Zhang (2011) can also be applied in the future.
@@ 3 Experiments
We conducted the experiments on the Fall Detection - Eldercare Robot (Elwaly and Abdellatif, 2023) dataset, with more details provided in Appendix A.1. We used a Vision-and-Language Transformer (ViLT) Kim et al. (2021) as the small model, chosen for its computational efficiency and relatively small size (less than 500 MB). For the large model, we utilized LLaVA-NeXT-Video-7B (Li et al., 2024), which is a well-performing open-source video-language model. Following the DMD process described in Section 2, we generated data where the inputs consist of the last hidden layer of the small model for each image input in the raw data, and the outputs indicate whether the small model and large model agree on their falling predictions. To minimize the computing on edge devices, we design a lightweight switcher model implemented as a Multi-Layer Perceptron (MLP) as an ad-hoc layer on top of the small model. The switcher model is trained on the DMD-generated data for the intelligent model switching. More implementation details are available in Appendix A.3.
@@ 3.1 Results
For performance comparison, we conducted experiments and included the results of three baselines. As reported in Table 1, reliance solely on the ViLT (Small model only) resulted in an F1 score of 58.2%, while LLaVA-NeXT-Video (Large model only) got 87.5%. In addition, we reproduced a popular uncertainty-based switcher model approach (Narayan et al., 2022) on our task. Specifically, we extracted the falling probability in the final output layer of ViLT and used that as the metric to determine whether to defer judgment to the larger model. This baseline got 76.1% when offloading 60% computing to the large model. Our system achieves 92.1%, which is 33.9%, 4.6%, and 16.0% higher than the traditional approach, large model-only approach and the uncertainty-based approach respectively.
@@
Small model only (Kim et al., 2021) 58.2 0 18.25 2.32
Large model only (Li et al., 2024) 87.5 100 663.1 190.73
Uncertainty-based system (Narayan et al., 2022) 76.1 60 386.91 113.04
@@ A.1 Datasets
We used the Fall Detection - Eldercare Robot dataset for training, validation and testing. It includes 1061 images, each having a data entry in YOLO v7 PyTorch format Wang et al. [2023]. The data entry includes a binary fall/no-fall classification label, the coordinates of the center of the bounding box containing the person, and the width and height of the bounding box. The creators of the dataset oriented pixel data automatically using EXIF-orientation stripping and stretched to resize it to 640x640. When exported, each image had a 50% probability of being laterally inverted and a one-in-three chance of undergoing one of the following operations: a 90-degree clockwise rotation, a 90-degree anti-clockwise rotation, or none at all. Of the 1061 images, we follow the commonly used split, i.e., 742 for training, 212 for validation, and 106 for testing.
@@ A.4 Threshold for Offload Decision
To explore how the threshold value influences the final testing performance, we further applied our "buckets" strategy to test data for both methods and analyzed the results. The corresponding results are shown in Figure 3. As a note, this analysis of accuracy against percent of data deferred was used extensively in [Narayan et al., 2022] and can be extended to any framework involving a switcher model.

### v2

@@ 1 Introduction
Recent advances in Artificial Intelligence (AI) have shown that increasing model size leads to substantial performance gains across various domains, from language models [3, 8] to video-language models [11, 18]. Large-scale models, with billions of parameters, consistently achieve state-of-the-art results on various benchmarks [19]. However, these models can be challenging to deploy when translating benchmark performance into real-world applications [13] due to their high runtime costs.
@@
One critical area where this limitation becomes evident is in action classification for elderly care [15]. A task such as predicting falls through non-intrusive monitoring [12, 5] is a critical yet complex task that benefits from large-scale video-language models [5]. Despite their high accuracy, these oversized models are not feasible for deployment on common elderly care appliances due to hardware constraints [1]. To address this, we propose a hybrid edge-cloud solution where small models run locally on edge devices while larger models deployed in the cloud are accessed only when necessary. However, the core technical challenge here remains: how to effectively align the knowledge and capabilities of the small and large models to ensure seamless collaboration between them. While Model or Knowledge Distillation techniques [7, 6] have been extensively explored, their primary goal is to compress the knowledge of large models into smaller ones. Yet, this process often falls short in collaborative systems where both small and large models have unique strengths. The smaller model, after distillation, may not fully exploit its distinct capabilities or optimally collaborate with its larger counterpart [14].
@@
To overcome these limitations, we introduce Dual-Model Distillation (DMD), a novel approach to enhance cooperation between small and large models. Unlike traditional distillation, which focuses on unidirectional knowledge transfer, DMD treats both models as teacher models to co-teach and generate synthetic data to train a student model (i.e., the switcher model). The DMD process eliminates the need for additional human-labeled data by leveraging the agreement or disagreement between the two teacher models’ zero-shot inferences to create training data. Through instruction-supervised finetuning [2], the switcher model learns to make optimal decisions about when to invoke the larger model. Experiments conducted on the dataset used in Fall Detection - Eldercare Robot datase [4] demonstrate the effectiveness of our approach. Surprisingly, our pipeline not only demands less computational overhead than just using the large model, resulting in 39.5% cost savings, but also outperforms the isolated large model by 4.6% in terms of F1 score. Furthermore, we anticipate that the DMD concept will provide substantial opportunities for future research on knowledge integration across multiple deep learning models.
@@ 2 Methodology
As shown in Figure 1, our hybrid edge-cloud solution introduces a companion switcher model using the hidden features of the input data extracted from the small model, which serves as the foundation of its reasoning capacity. With the help of the switcher model, our edge model can not only predict the falling probability but also accurately estimate the probability that the large model’s result aligns with its own. By incorporating this switcher model, the edge model deliberates when to invoke the cloud model, thereby keeping system expenses under a pre-defined budget. In our case, a threshold of alignment probability, which yields optimal performance in predicting the falling probability during training, determines when to offload data to the large model during inference time. Regarding more complex decision-making scenarios, we expect that a more sophisticated utility-based threshold design [17] can also be applied in the future.
@@ 3 Experiments
We conducted the experiments on the Fall Detection - Eldercare Robot [4] dataset, with more details provided in Appendix A.1. We used a Vision-and-Language Transformer (ViLT) [9] as the small model, chosen for its computational efficiency and relatively small size (less than 500 MB). For the large model, we utilized LLaVA-NeXT-Video-7B [10], which is a well-performing open-source video-language model. Following the DMD process described in Section 2, we generated data where the inputs consist of the last hidden layer of the small model for each image input in the raw data, and the outputs indicate whether the small model and large model agree on their falling predictions. To minimize the computing on edge devices, we design a lightweight switcher model implemented as a Multi-Layer Perceptron (MLP) as an ad-hoc layer on top of the small model. The switcher model is trained on the DMD-generated data for the intelligent model switching. More implementation details are available in Appendix A.3.
@@ 3.1 Results
For performance comparison, we conducted experiments and included the results of three baselines. As reported in Table 1, reliance solely on the ViLT (Small model only) resulted in an F1 score of 58.2%, while LLaVA-NeXT-Video (Large model only) got 87.5%. In addition, we reproduced a popular uncertainty-based switcher model approach [14] on our task. Specifically, we extracted the falling probability in the final output layer of ViLT and used that as the metric to determine whether to defer judgment to the larger model. This baseline got 76.1% when offloading 60% computing to the large model. Our system achieves 92.1%, which is 33.9%, 4.6%, and 16.0% higher than the traditional approach, large model-only approach and the uncertainty-based approach respectively.
@@
Small model only [9] 58.2 0 18.25 2.32
Large model only [10] 87.5 100 663.1 190.73
Uncertainty-based system [14] 76.1 60 386.91 113.04
@@ 4 Conclusion and Future Works
Acknowledgments and Disclosure of Funding
We would like to thank Evan Li and Ryan Tellado for their insightful feedback and contributions to the development of this paper.
@@ A.1 Datasets
We used the Fall Detection - Eldercare Robot dataset for training, validation and testing. It includes 1061 images, each having a data entry in YOLO v7 PyTorch format [16]. The data entry includes a binary fall/no-fall classification label, the coordinates of the center of the bounding box containing the person, and the width and height of the bounding box. The creators of the dataset oriented pixel data automatically using EXIF-orientation stripping and stretched to resize it to 640x640. When exported, each image had a 50% probability of being laterally inverted and a one-in-three chance of undergoing one of the following operations: a 90-degree clockwise rotation, a 90-degree anti-clockwise rotation, or none at all. Of the 1061 images, we follow the commonly used split, i.e., 742 for training, 212 for validation, and 106 for testing.
@@ A.4 Threshold for Offload Decision
To explore how the threshold value influences the final testing performance, we further applied our "buckets" strategy to test data for both methods and analyzed the results. The corresponding results are shown in Figure 3. As a note, this analysis of accuracy against percent of data deferred was used extensively in [14] and can be extended to any framework involving a switcher model.